OpenStudy (this):

helpppppp plllzzzz find the zeros of the function y=x^2-64 a,(8,0)and(8,0) b,(8,0)and (-8,0) c,(-8,0)and (-8,0) d,(4,0)and(0,0)

OpenStudy (anonymous):

x^2-64=x^2-8^2 does that help

OpenStudy (anonymous):

difference of two squares

hero (hero):

y = x^2 - 64 Set y = 0, then solve for x: 0 = x^2 - 64 64 = x^2 Take square root, both sides: ±8 = x Thus: (8,0) and (-8,0)

OpenStudy (anonymous):

\[a^2-b^2=(a-b)(a+b)\]